OpenStudy (anonymous):

a regular octagon with side of length 8 and a apothem of length 9.66 has a area of?

OpenStudy (zpupster):

A= 1/2*a*n*s a= apothem n= number of sides s= side length A = 1/2(9.66)(8)(8)

OpenStudy (anonymous):

Thanks @zpupster